OpenStudy (anonymous):

How do I convert the rectangular coordinates (5,7) to polar coordinates?

OpenStudy (anonymous):

What are rectangular and polar coordinates?

OpenStudy (anonymous):

x = r cos(theta) y = r sin(theta)

OpenStudy (anonymous):

hm?

OpenStudy (anonymous):

(x,y) x = r cos(theta) y = r sin(theta)

OpenStudy (anonymous):

x^2 + y^2 = r^2

OpenStudy (anonymous):

what is r and theta

OpenStudy (anonymous):

theta is the angle you should have learned it by now if you are doing polar coordinates

OpenStudy (anonymous):

Well, I'm a freshman and this is just on an academic team test, so I don't know any of it

OpenStudy (anonymous):



OpenStudy (anonymous):

ah, that's nice

OpenStudy (anonymous):

How the hell did you draw that perfectly? mine was completely disgusting...

OpenStudy (anonymous):

so rectangular coordinates just means cartesian coordinates?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, what next?

OpenStudy (anonymous):

theta = arctan(y/x) r= x/cos(theta)

OpenStudy (anonymous):

what is arctan?

OpenStudy (anonymous):

inverse operation of tangent

OpenStudy (anonymous):

tan^-1 like on my calculator?

OpenStudy (anonymous):

[INV] [TAN] on your calculator

OpenStudy (anonymous):

probably.

OpenStudy (anonymous):

Dont use the calculator..learn it

OpenStudy (anonymous):

you using Google's calculator? Then yes it's tan^-1

OpenStudy (anonymous):

no, my ti-84

OpenStudy (anonymous):

2nd tan = tan^-1

OpenStudy (anonymous):

I think that's labeled tan^-1 also... don't remember for sure...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so would doing arctan(7/5) give me the actual angle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you explain to me why I use arctan please?

OpenStudy (anonymous):

oh wait, I think i see

OpenStudy (anonymous):

tangent is opp/adj which is 7/5 which gives you the actual tangent, and then convert it to the angle?

OpenStudy (anonymous):

tan(theta) = opp/adj = y/x tan^-1(tan(theta)) = tan^-1 (y/x) theta = tan^-1(y/x)

OpenStudy (anonymous):

yes, exactly.

OpenStudy (anonymous):

sweet

OpenStudy (anonymous):

and then we have our angle which is about 54.46 degrees

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we need to find r which is sqrt(25+49)?

OpenStudy (anonymous):

you can...

OpenStudy (anonymous):

or r= x/cos(theta) or r= y/sin(theta)

OpenStudy (anonymous):

can you explain to me why the 3 are the same thing?

OpenStudy (anonymous):

better to use the sqrt. it's valid for all quadrants, since r is usually reckoned positive (except in some quirky cases). ...since this is 1st quadrant you can use the 'shortcuts' I gave...

OpenStudy (anonymous):

ok, would the way we found theta always work or are their exceptions too?

OpenStudy (anonymous):

there*

OpenStudy (anonymous):

Sort of... calculators give values for theta (when finding it by arctan) that are 1st or 4th quadrants...

OpenStudy (anonymous):